OpenStudy (toxicsugar22):

acetic acid =4.76, ka=_________M

OpenStudy (radar):

I'm cluless!

OpenStudy (sleepyjess):

Sorry i have no idea

OpenStudy (sleepyjess):

@ganeshie8 maybe?

OpenStudy (anonymous):

I think it is... pH = 4.760 + log 0.00379/ 0.00621 = 4.55 Not really sure. It may be wrong though...

OpenStudy (toxicsugar22):

@Sepeario

OpenStudy (toxicsugar22):

@surjithayer @shrutipande9 @sammixboo @ShadowLegendX @sammie2017 @Donblue22 @Dark_Riddles @rishavraj @anonymous_user @april115 @Awesome781 @Frosty_Flow @tester97 @thadyoung @iambatman @ikram002p @ParthKohli @genny7 @grray7 @ganeshie8 @Haseeb96 @happyyanee2 @hba @Jumperman @lissiewalkin @lilfayfay @Lrki @zubhanwc3 @Cydney_Morgan @bibby @bluebrandon @bestie @nirmalnema @navina @missylulu @Mclovin21 @mony01

OpenStudy (jumperman):

*YAWN* Huh? You woke me from my hibernation? Don't you see that I'm a wolf?

OpenStudy (haseeb96):

Sorry.

OpenStudy (toxicsugar22):

Does anyone know how to do this

OpenStudy (anonymous):

You just woke up a wolf, man IDEK how to do this, google Helps.

OpenStudy (toxicsugar22):

acetic acid pKa=4.76, ka=_________M

OpenStudy (jumperman):

Don't ask a wolf! Wolves only know about how to sleep, eat, hunt, howl, and.... other stuff. Not this, it's not important to our way of life.

OpenStudy (nirmalnema):

You know that the pH = 2.88 pH = -log [H+] using log laws: [H+] = 10^-pH [H+] = 10^-2.88 [H+] = 1.318*10^-3 You know that : For acetic acid use HAc Ka = [H+] *[Ac]/ [HAc] you know [H+] = [Ac-] = 1.318*10^-3 . So [H+] * [Ac-] = [H+]² and [HAc] = 0.1 Substitute: Ka = (1.318*10^-3)²/0.1 Ka = 1.74*10^-6 / 0.1 Ka = 1.74*10^-5 pKa = -log Ka pKa = -log ( 1.74*10^-5) pKa = 4.76

OpenStudy (nirmalnema):

@toxicsugar22 do u get that ??